# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v2

@@ 5.6 Deep Learning-Based Missing Data Imputation
Deep learning-based methods for missing data imputation (KHAN2022278) can be categorized into three groups: autoencoder-based, deep matrix factorization, and deep generative model-based methods. These categories are detailed in the subsequent sections.
@@ 5.6.1 Autoencoder-Based Methods
An autoencoder (AE) (hinton2006reducing) is a type of artificial neural network (lecun2015deep) used for unsupervised learning of efficient codings. The goal of an AE is to learn a representation (encoding) for a set of data, typically for the purpose of dimensionality reduction or feature learning. Given a dataset with $n$ data points of $m$-dimension, a standard stacked or deep AE is composed of an encoder network $h_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ and a decoder network $g_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, which is often trained by solving the following optimization
@@
$\mathbf{z}=h_{\theta}(\mathbf{x})$ is usually called the representation of $\mathbf{x}$. Note that for tabular data, $h_{\theta}$ and $g_{\phi}$ are usually multilayer perceptrons, e.g., $h_{\theta}(\mathbf{x})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, where $\theta=\{\mathbf{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}}\}$ are the set of parameters to optimize and $\sigma$ denote the activation function. For image data, $h_{\theta}$ and $g_{\phi}$ are usually based on convolutional layers. Variants of the standard autoencoder include the denoising autoencoder (DAE) (vincent2010stacked), the sparse autoencoder (ng2011sparse), and the variational autoencoder (VAE) (kingma2013auto).
vincent2008 used DAE to recover missing values of data, but it requires complete data during training. Applying AE or DAE to incomplete data directly, where the reconstruction errors for the observed values are minimized, works in practice (beaulieu2017missing; ryu2020denoising):
@@
where $\bm{\omega}_{i}\in\{0,1\}^{m}$ is the missing value indicator for $\mathbf{x}_{i}$ and the missing values in the input are often temporarily filled with zeros. However, the input bias caused by the missing values could be harmful. FAN2017540 proposed a deep learning based matrix completion method, which optimizes the missing values and the autoencoder parameters simultaneously and hence can reduce the input bias caused by the missing values. gondara2018mida proposed a multiple imputation model based on DAE. pereira2020reviewing wrote a review paper of AE and its variants based missing data imputation methods. fan2024neuronenhanced proposed a neuron-enhanced autoencoder for missing data imputation, where the activation function of the output layer is learned adaptively by a neural network to model the complex response function in real data. du2024remasker proposed a simple yet effective imputation method called ReMasker. In ReMasker, besides the missing values (i.e., naturally masked), another set of values is randomly re-masked, and the autoencoder is optimized by reconstructing the re-masked set. ReMasker performs on par with or outperforms many complex and strong competitors in terms of both imputation fidelity and utility under various missingness settings.
@@ 5.6.2 Deep Matrix Factorization-Based Methods
AE-based imputation models need to learn to reduce the negative impact of the missing values in the input of the encoder, which can be difficult in practice. To address the limitation, (fan2018matrix) proposed deep matrix factorization (DMF) based matrix completion. Instead of a linear and shallow factorization like (11), they solve the following deep factorization problem
@@
where the bias terms of the neural network are omitted for simplicity and R{R} denotes some regularization on the latent matrix $\mathbf{Z}\in\mathbb{R}^{d\times n}$ and the weight matrices. One limitation of this model is that the latent dimension $d$ is difficult to determine in real-world applications. There have been a few extensions of DMF-based missing data imputation (zhang2019deep; li2022regularised; li2023h2tf; ledent2024generalization; ji2025deep).
@@ 5.6.3 Deep Generative Model-Based Methods
Deep generative models are a class of neural network architectures designed to learn the underlying distribution of a dataset and generate new data samples that resemble the training data. These models are particularly useful for tasks such as image synthesis, text generation, and data augmentation. They have gained significant attention in recent years due to their ability to produce high-quality and realistic data. Key types of deep generative models include Variational Autoencoders (VAEs) (kingma2013auto), Generative Adversarial Networks (GANs) (goodfellow2014generative; gulrajani2017improved), Normalizing Flows (NFs) (dinh2014nice; kingma2016improved), Autoregressive Models (van2016pixel; salimans2017pixelcnn++), and Diffusion models (DMs) (sohl2015deep; ho2020denoising).
@@
The basic VAE (kingma2013auto) solves the following problem
@@
The encoder $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ approximates the posterior distribution over latent variables $\mathbf{z}\in\mathbb{R}^{d}$ given input $\mathbf{x}\in\mathbb{R}^{m}$ and outputs the parameters of a prior distribution (e.g., the mean and variance of a Gaussian). The decoder $p_{\theta}(\mathbf{x}\mid\mathbf{z})$ reconstructs the input $\mathbf{x}$ from latent variable $\mathbf{z}$, since maximizing $\log p_{\theta}(\mathbf{x}\mid\mathbf{z})$ is equivalent to minimizing the reconstruction error. The KL Divergence $D_{\mathrm{KL}}$ regularizes the latent space by pushing $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ toward a prior $p(\mathbf{z})$ (e.g., standard Gaussian $\mathcal{N}(\bm{0},\mathbf{I})$). $\mathcal{L}(\theta,\phi)$ is a lower bound of the log-likelihood $\log p_{\theta}(\mathbf{x})$. Once the model is trained, new data points can be generated by sampling a latent vector $\mathbf{z}$ from the prior $p(\mathbf{z})$ and passing it through the decoder. The core VAE framework has inspired several variants (burda2015importance; higgins2017beta; pmlr-v97-mathieu19a).
The MIWAE method proposed by (mattei2019miwae) adapts the importance-weighted autoencoder (IWAE) (burda2015importance) to a missing data imputation model, where IWAE shares the same architecture as VAE but has a different training objective and a tighter lower bound of the log-likelihood. Specifically, letting $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ be the observed and missing parts of $\mathbf{x}_{i}$, respectively, MIWAE solves
@@
There are more VAE-based imputation methods. For example, nazabal2020handling proposed HI-VAE for heterogeneous (mixed continuous and discrete) data imputation. collier2020vaes proposed a VAE-based imputation method that explicitly models missing values via a latent corruption process, enabling principled handling of both MCAR and MNAR data by conditioning the encoder and decoder on missingness indicators. fortuin2020gp proposed a method called GP-VAE for time series imputation.
@@
The basic GAN (goodfellow2014generative) solves the following problem
@@
where $G$ and $D$ are the generator and discriminator, respectively. $G$ is a neural network that learns to generate fake data from random noise $\mathbf{z}$, usually drawn from a Gaussian distribution. It tries to minimize $V(D,G)$ (fool the discriminator). $D$, usually a binary classifier, learns to distinguish real data ($\mathbf{x}$, labeled as $1$) from fake data ($G(\mathbf{z})$, labeled as 0). It tries to maximize $V(D,G)$. Representative variants of GAN include cycle-GAN (zhu2017unpaired), Wasserstein GAN (arjovsky2017wasserstein), etc.
A wide range of GAN-based methods have been developed for missing data imputation (shang2017vigan; yoon2018gain; luo2018multivariate; li2019misgan; yang2020adversarial; gupta2020time; zhang2021missing; dong2021generative; xia2021recovering; kazemi2021igani; wang2022sta; qin2023imputegan; kang2023cm; wang2024time; wang2025generative). shahbazian2023generative provided a review of GAN-based imputation methods, including a numerical performance evaluation of several models. A common paradigm in these approaches is to use the generator as an imputation network, which takes as input either a random latent vector $\mathbf{z}$, the incomplete data $\mathbf{x}$, or a combination of both with additional contextual information. Taking the GAIN proposed by (yoon2018gain) as an example, let $\tilde{\mathbf{x}}$, $\bm{\omega}$, and $\mathbf{z}$ be the incomplete data vector, binary mask vector, and random noise, respectively. The generator in GAIN is designed to impute missing data, i.e., $\hat{\mathbf{x}}=G(\tilde{\mathbf{x}},\bm{\omega},(\mathbf{1}-\bm{\omega})\odot\mathbf{z})$, while the discriminator is tasked with determining whether a particular element is observed or imputed, i.e., $\hat{\bm{\omega}}=D(\bar{\mathbf{x}},\mathbf{h})$, where $\bar{\mathbf{x}}=\bm{\omega}\odot\tilde{\mathbf{x}}+(\mathbf{1}-\bm{\omega})\odot\hat{\mathbf{x}}$ and $\mathbf{h}$ is a hint vector generated by letting one or more randomly selected entries of $\bm{\omega}$ be $0.5$. The loss functions for the discriminator and generator are as follows:
@@
Normalizing flows (dinh2014nice; kingma2016improved), which transform a simple base distribution $q(\mathbf{z})$ (e.g., a standard Gaussian) into a complex target distribution $p(\mathbf{x})$ through a bijective function $\mathbf{x}=g(\mathbf{z})$:
@@
where $\lambda>0$ is a hyperparameter. Please note that the notation used here differs from that in (richardson2020mcflow). This change was made to maintain consistency with the mathematical formulations of the previously introduced methods.
@@
Diffusion models (sohl2015deep; ho2020denoising; song2020score; yang2023diffusion) are generative models that learn to generate data (e.g., images, audio) by gradually denoising random noise into structured samples. They consist of two key processes. The first is the forward (diffusion) process, which gradually corrupts the training data with Gaussian noise over many steps, transforming it into pure noise, i.e., at each step $t$,
@@
Since diffusion models are often more effective than other generative models, such as VAEs and GANs, in generating high-quality new samples, a few researchers have attempted to extend them to missing data imputation. For instance, tashiro2021csdi proposed CSDI, a diffusion model-based time series imputation method, by modeling the following distribution
@@
Inspired by CSDI, zheng2022diffusion applied diffusion models to tabular data imputation. ouyang2023missdiff proposed MissDiff for tabular data imputation. chen2024rethinking introduced negative entropy-regularized Wasserstein gradient flow for diffusion model-based imputation. jolicoeur2024generating introduced an approach for generating and imputing mixed-type (continuous and categorical) tabular data utilizing score-based diffusion and conditional flow matching. liu2024self demonstrated that self-supervised pretraining enhances diffusion models for tabular data imputation by leveraging unlabeled data to learn richer feature representations before fine-tuning on missing-value tasks. zhang2025diffputer proposed DiffPuter, which iteratively trains a diffusion model to learn the joint distribution of missing and observed data and performs an accurate conditional sampling to update the missing values using a tailored reversed sampling strategy. ahamed2025refidiff proposed RefiDiff that combines local machine learning predictions with a novel Mamba-based denoising network capturing interrelationships among distant features and samples, which improved the imputation performance of diffusion models.
@@ 5.6.4 Graph Neural Network-Based Methods
Representative architectures of GNNs include graph convolutional network (GCN) (kipf2017semi), graph attention network GAT) (velivckovic2017graph), graph isomorphism network (GIN) (xu2019powerful), and graph transformer (GT) (yun2019graph). The hidden layers of a GCN can be formulated as
@@
Several studies have explored using Graph Neural Networks (GNNs) for missing data imputation (zhang2019inductive; you2020handling; spinelli2020missing; telyatnikov2023egg). A key challenge, however, is that GNNs typically require a pre-defined similarity graph across the entire dataset, which is difficult to construct from incomplete data. zhang2019inductive proposed a GNN-based matrix completion method for collaborative filtering. you2020handling proposed GRAPE, a graph-based framework for feature imputation as well as label prediction. GRAPE tackles the missing data problem using a graph representation, where the observations and features are viewed as two types of nodes in a bipartite graph, and the observed feature values as edges. Specifically, in GRAPE, for the incomplete data matrix $\tilde{\mathbf{X}}$ (with mask matrix $\mathbf{M}$) containing $n$ observations and $d$ attributes, a bipartite graph $G=(V_{n}\cup V_{d},E)$ is constructed, where $V_{n}=\{u_{1},u_{2},\ldots,u_{n}\}$ denotes the set of nodes corresponding to the $n$ observations, $V_{d}=\{v_{1},v_{2},\ldots,v_{d}\}$ denotes the set of nodes corresponding to the $d$ attributes, and $E=\left\{\left(u_{i},v_{j},e_{u_{i},v_{j}}\right):u_{i}\in V_{n},v_{j}\in V_{d},e_{u_{i},v_{j}}=\tilde{x}_{ij},M_{ij}=1\right\}$ denotes the edge set. The prediction for the missing values in $\tilde{\mathbf{X}}$ is given by
@@
where $f_{\theta}$ is a GNN such as GraphSAGE (hamilton2017inductive). $\mathbf{H}_{n}\in\{1\}^{n\times d}$ is the input embedding matrix for the $n$ sample nodes and $\mathbf{H}_{d}=\mathbf{I}_{d}$ is the input embedding matrix for the $d$ attribute nodes; they are updated in the subsequent layers via neighbor aggregation over $G$. The parameters $\theta$ of the GNN are then optimized using a loss function such as $\sum_{(i,j):M_{ij}=1}(\tilde{x}_{ij}-\hat{x}_{ij})^{2}$.
More recently, telyatnikov2023egg proposed EGG-GAE, a GCN-based method for tabular data that adaptively constructs the adjacency matrix from a neural network’s representations of the incomplete data. In a different domain, li2024stmcdi combined a GNN with a diffusion model to impute missing values in spatial transcriptomics data. More methods of GNN-based imputation can be found in (huang2022graph; wu2023overview; eskandari2024gn2di; zhang2025enhancing).

### v3

@@ 5.6 Deep Learning-Based Missing Data Imputation
Deep learning-based methods for missing data imputation (Khan et al., 2022) can be categorized into three groups: autoencoder-based, deep matrix factorization, and deep generative model-based methods. These categories are detailed in the subsequent sections.
@@ 5.6.1 Autoencoder-Based Methods
An autoencoder (AE) (Hinton and Salakhutdinov, 2006) is a type of artificial neural network (LeCun et al., 2015) used for unsupervised learning of efficient codings. The goal of an AE is to learn a representation (encoding) for a set of data, typically for the purpose of dimensionality reduction or feature learning. Given a dataset with $n$ data points of $m$-dimension, a standard stacked or deep AE is composed of an encoder network $h_{\theta}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ and a decoder network $g_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, which is often trained by solving the following optimization
@@
$\mathbf{z}=h_{\theta}(\mathbf{x})$ is usually called the representation of $\mathbf{x}$. Note that for tabular data, $h_{\theta}$ and $g_{\phi}$ are usually multilayer perceptrons, e.g., $h_{\theta}(\mathbf{x})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, where $\theta=\{\mathbf{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}}\}$ are the set of parameters to optimize and $\sigma$ denote the activation function. For image data, $h_{\theta}$ and $g_{\phi}$ are usually based on convolutional layers. Variants of the standard autoencoder include the denoising autoencoder (DAE) (Vincent et al., 2010), the sparse autoencoder (Ng and others, 2011), and the variational autoencoder (VAE) (Kingma and Welling, 2013).
Vincent et al. (2008) used DAE to recover missing values of data, but it requires complete data during training. Applying AE or DAE to incomplete data directly, where the reconstruction errors for the observed values are minimized, works in practice (Beaulieu-Jones et al., 2017; Ryu et al., 2020):
@@
where $\bm{\omega}_{i}\in\{0,1\}^{m}$ is the missing value indicator for $\mathbf{x}_{i}$ and the missing values in the input are often temporarily filled with zeros. Figure 13 shows an example of AE-based imputation. However, the input bias caused by the missing values could be harmful. Fan and Chow (2017c) proposed a deep learning based matrix completion method, which optimizes the missing values and the autoencoder parameters simultaneously and hence can reduce the input bias caused by the missing values. Gondara and Wang (2018) proposed a multiple imputation model based on DAE. Pereira et al. (2020) wrote a review paper of AE and its variants based missing data imputation methods. Fan et al. (2024) proposed a neuron-enhanced autoencoder for missing data imputation, where the activation function of the output layer is learned adaptively by a neural network to model the complex response function in real data. Du et al. (2024a) proposed a simple yet effective imputation method called ReMasker. In ReMasker, besides the missing values (i.e., naturally masked), another set of values is randomly re-masked, and the autoencoder is optimized by reconstructing the re-masked set. ReMasker performs on par with or outperforms many complex and strong competitors in terms of both imputation fidelity and utility under various missingness settings.
Figure 13: A simple example of AE-based imputation. The rows of the matrix correspond to attributes. The neural network has 3 hidden layers. The missing values in the input layer are set to zero.
@@ 5.6.2 Deep Matrix Factorization-Based Methods
AE-based imputation models need to learn to reduce the negative impact of the missing values in the input of the encoder, which can be difficult in practice. To address the limitation, (Fan and Cheng, 2018) proposed deep matrix factorization (DMF) based matrix completion. Instead of a linear and shallow factorization like (11), they solve the following deep factorization problem
@@
where the bias terms of the neural network are omitted for simplicity and R{R} denotes some regularization on the latent matrix $\mathbf{Z}\in\mathbb{R}^{d\times n}$ and the weight matrices. One limitation of this model is that the latent dimension $d$ is difficult to determine in real-world applications. There have been a few extensions of DMF-based missing data imputation (Zhang et al., 2019b; Li et al., 2022, 2023a; Ledent and Alves, 2024; Ji et al., 2025).
@@ 5.6.3 Deep Generative Model-Based Methods
Deep generative models are a class of neural network architectures designed to learn the underlying distribution of a dataset and generate new data samples that resemble the training data. These models are particularly useful for tasks such as image synthesis, text generation, and data augmentation. They have gained significant attention in recent years due to their ability to produce high-quality and realistic data. Key types of deep generative models include Variational Autoencoders (VAEs) (Kingma and Welling, 2013), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Gulrajani et al., 2017), Normalizing Flows (NFs) (Dinh et al., 2014; Kingma et al., 2016), Autoregressive Models (Van Den Oord et al., 2016; Salimans et al., 2017), and Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020).
@@
The basic VAE (Kingma and Welling, 2013) solves the following problem
@@
The encoder $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ approximates the posterior distribution over latent variables $\mathbf{z}\in\mathbb{R}^{d}$ given input $\mathbf{x}\in\mathbb{R}^{m}$ and outputs the parameters of a prior distribution (e.g., the mean and variance of a Gaussian). The decoder $p_{\theta}(\mathbf{x}\mid\mathbf{z})$ reconstructs the input $\mathbf{x}$ from latent variable $\mathbf{z}$, since maximizing $\log p_{\theta}(\mathbf{x}\mid\mathbf{z})$ is equivalent to minimizing the reconstruction error. The KL Divergence $D_{\mathrm{KL}}$ regularizes the latent space by pushing $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ toward a prior $p(\mathbf{z})$ (e.g., standard Gaussian $\mathcal{N}(\bm{0},\mathbf{I})$). $\mathcal{L}(\theta,\phi)$ is a lower bound of the log-likelihood $\log p_{\theta}(\mathbf{x})$. Once the model is trained, new data points can be generated by sampling a latent vector $\mathbf{z}$ from the prior $p(\mathbf{z})$ and passing it through the decoder. The core VAE framework has inspired several variants (Burda et al., 2015; Higgins et al., 2017; Mathieu et al., 2019).
The MIWAE method proposed by (Mattei and Frellsen, 2019) adapts the importance-weighted autoencoder (IWAE) (Burda et al., 2015) to a missing data imputation model, where IWAE shares the same architecture as VAE but has a different training objective and a tighter lower bound of the log-likelihood. Specifically, letting $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ be the observed and missing parts of $\mathbf{x}_{i}$, respectively, MIWAE solves
@@
There are more VAE-based imputation methods. For example, Nazabal et al. (2020) proposed HI-VAE for heterogeneous (mixed continuous and discrete) data imputation. Collier et al. (2020) proposed a VAE-based imputation method that explicitly models missing values via a latent corruption process, enabling principled handling of both MCAR and MNAR data by conditioning the encoder and decoder on missingness indicators. Fortuin et al. (2020) proposed a method called GP-VAE for time series imputation.
@@
The basic GAN (Goodfellow et al., 2014) solves the following problem
@@
where $G$ and $D$ are the generator and discriminator, respectively. $G$ is a neural network that learns to generate fake data from random noise $\mathbf{z}$, usually drawn from a Gaussian distribution. It tries to minimize $V(D,G)$ (fool the discriminator). $D$, usually a binary classifier, learns to distinguish real data ($\mathbf{x}$, labeled as $1$) from fake data ($G(\mathbf{z})$, labeled as 0). It tries to maximize $V(D,G)$. Representative variants of GAN include cycle-GAN (Zhu et al., 2017), Wasserstein GAN (Arjovsky et al., 2017), etc.
A wide range of GAN-based methods have been developed for missing data imputation (Shang et al., 2017; Yoon et al., 2018; Luo et al., 2018; Li et al., 2019; Yang et al., 2020; Gupta and Beheshti, 2020; Zhang et al., 2021b; Dong et al., 2021; Xia et al., 2021; Kazemi and Meidani, 2021; Wang et al., 2022b; Qin and Wang, 2023; Kang et al., 2023; Wang et al., 2024d, 2025). Shahbazian and Greco (2023) provided a review of GAN-based imputation methods, including a numerical performance evaluation of several models. A common paradigm in these approaches is to use the generator as an imputation network, which takes as input either a random latent vector $\mathbf{z}$, the incomplete data $\mathbf{x}$, or a combination of both with additional contextual information. Taking the GAIN proposed by (Yoon et al., 2018) as an example, let $\tilde{\mathbf{x}}$, $\bm{\omega}$, and $\mathbf{z}$ be the incomplete data vector, binary mask vector, and random noise, respectively. The generator in GAIN is designed to impute missing data, i.e., $\hat{\mathbf{x}}=G(\tilde{\mathbf{x}},\bm{\omega},(\mathbf{1}-\bm{\omega})\odot\mathbf{z})$, while the discriminator is tasked with determining whether a particular element is observed or imputed, i.e., $\hat{\bm{\omega}}=D(\bar{\mathbf{x}},\mathbf{h})$, where $\bar{\mathbf{x}}=\bm{\omega}\odot\tilde{\mathbf{x}}+(\mathbf{1}-\bm{\omega})\odot\hat{\mathbf{x}}$ and $\mathbf{h}$ is a hint vector generated by letting one or more randomly selected entries of $\bm{\omega}$ be $0.5$. The loss functions for the discriminator and generator are as follows:
@@
Normalizing flows (Dinh et al., 2014; Kingma et al., 2016), which transform a simple base distribution $q(\mathbf{z})$ (e.g., a standard Gaussian) into a complex target distribution $p(\mathbf{x})$ through a bijective function $\mathbf{x}=g(\mathbf{z})$:
@@
where $\lambda>0$ is a hyperparameter. Please note that the notation used here differs from that in (Richardson et al., 2020). This change was made to maintain consistency with the mathematical formulations of the previously introduced methods.
@@
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Yang et al., 2023) are generative models that learn to generate data (e.g., images, audio) by gradually denoising random noise into structured samples. They consist of two key processes. The first is the forward (diffusion) process, which gradually corrupts the training data with Gaussian noise over many steps, transforming it into pure noise, i.e., at each step $t$,
@@
Since diffusion models are often more effective than other generative models, such as VAEs and GANs, in generating high-quality new samples, a few researchers have attempted to extend them to missing data imputation. For instance, Tashiro et al. (2021) proposed CSDI, a diffusion model-based time series imputation method, by modeling the following distribution
@@
Inspired by CSDI, Zheng and Charoenphakdee (2022) applied diffusion models to tabular data imputation. Ouyang et al. (2023) proposed MissDiff for tabular data imputation. Chen et al. (2024b) introduced negative entropy-regularized Wasserstein gradient flow for diffusion model-based imputation. Jolicoeur-Martineau et al. (2024) introduced an approach for generating and imputing mixed-type (continuous and categorical) tabular data utilizing score-based diffusion and conditional flow matching. Liu et al. (2024) demonstrated that self-supervised pretraining enhances diffusion models for tabular data imputation by leveraging unlabeled data to learn richer feature representations before fine-tuning on missing-value tasks. Zhang et al. (2025a) proposed DiffPuter, which iteratively trains a diffusion model to learn the joint distribution of missing and observed data and performs an accurate conditional sampling to update the missing values using a tailored reversed sampling strategy. Ahamed et al. (2025) proposed RefiDiff that combines local machine learning predictions with a novel Mamba-based denoising network capturing interrelationships among distant features and samples, which improved the imputation performance of diffusion models.
@@ 5.6.4 Graph Neural Network-Based Methods
Representative architectures of GNNs include graph convolutional network (GCN) (Kipf and Welling, 2017), graph attention network GAT) (Veličković et al., 2017), graph isomorphism network (GIN) (Xu et al., 2019a), and graph transformer (GT) (Yun et al., 2019). The hidden layers of a GCN can be formulated as
@@
Several studies have explored using Graph Neural Networks (GNNs) for missing data imputation (Zhang and Chen, 2020; You et al., 2020; Spinelli et al., 2020; Telyatnikov and Scardapane, 2023). A key challenge, however, is that GNNs typically require a pre-defined similarity graph across the entire dataset, which is difficult to construct from incomplete data. Zhang and Chen (2020) proposed a GNN-based matrix completion method for collaborative filtering. You et al. (2020) proposed GRAPE, a graph-based framework for feature imputation as well as label prediction. GRAPE tackles the missing data problem using a graph representation, where the observations and features are viewed as two types of nodes in a bipartite graph, and the observed feature values as edges. Specifically, in GRAPE, for the incomplete data matrix $\tilde{\mathbf{X}}$ (with mask matrix $\mathbf{M}$) containing $n$ observations and $d$ attributes, a bipartite graph $G=(V_{n}\cup V_{d},E)$ is constructed, where $V_{n}=\{u_{1},u_{2},\ldots,u_{n}\}$ denotes the set of nodes corresponding to the $n$ observations, $V_{d}=\{v_{1},v_{2},\ldots,v_{d}\}$ denotes the set of nodes corresponding to the $d$ attributes, and $E=\left\{\left(u_{i},v_{j},e_{u_{i},v_{j}}\right):u_{i}\in V_{n},v_{j}\in V_{d},e_{u_{i},v_{j}}=\tilde{x}_{ij},M_{ij}=1\right\}$ denotes the edge set. The prediction for the missing values in $\tilde{\mathbf{X}}$ is given by
@@
where $f_{\theta}$ is a GNN such as GraphSAGE (Hamilton et al., 2017). $\mathbf{H}_{n}\in\{1\}^{n\times d}$ is the input embedding matrix for the $n$ sample nodes and $\mathbf{H}_{d}=\mathbf{I}_{d}$ is the input embedding matrix for the $d$ attribute nodes; they are updated in the subsequent layers via neighbor aggregation over $G$. The parameters $\theta$ of the GNN are then optimized using a loss function such as $\sum_{(i,j):M_{ij}=1}(\tilde{x}_{ij}-\hat{x}_{ij})^{2}$.
More recently, Telyatnikov and Scardapane (2023) proposed EGG-GAE, a GCN-based method for tabular data that adaptively constructs the adjacency matrix from a neural network’s representations of the incomplete data. In a different domain, Li et al. (2024b) combined a GNN with a diffusion model to impute missing values in spatial transcriptomics data. More methods of GNN-based imputation can be found in (Huang et al., 2022; Wu et al., 2023; Eskandari et al., 2024; Zhang et al., 2025d).